OpenStudy (owlet):

A gaseous reaction occurs at a constant pressure of 50.4 bar and releases 67.0 kJ of heat. Before the reaction, the volume of the system was 9.00 L . After the reaction, the volume of the system was 2.20 L . Calculate the total internal energy change, ΔU, in kilojoules.

OpenStudy (owlet):

So I know.. ΔU = q + W and W = - PΔV q=67.0 kJ ΔV = 2.20-9L= -6.8L P= 50.4 bar I also know 1 L•bar = 100 J Substituting all my given values: ΔU = q + W ΔU = -67.0 kJ + (-50.4 bar)(-6.8 L ) ΔU = -67.0 kJ + 342.72 L•bar ΔU = -67.0 kJ + 34.272 kJ ΔU = -32.728 kJ AM I RIGHT?

OpenStudy (owlet):

@aaronq

OpenStudy (aaronq):

yep! you got the signs and everything right, it looks right (as long as you didn't make a mistake multiplying or something)

OpenStudy (owlet):

okay thank you! I just wanted to make sure that it is right :)

OpenStudy (aaronq):

no problem! is this from a textbook? they usually have answers at the back, or you can download the solution manual